The Heuser Company's currently outstanding bonds have a 8% coupon and a 13% yield to maturity. Heuser believes it could issue new bonds at par that would provide a similar yield to maturity. If its marginal tax rate is 40%, what is Heuser's after-tax cost of debt? Round your answer to two decimal places.

Tunney Industries can issue perpetual preferred stock at a price of $71.00 a share. The stock would pay a constant annual dividend of $6.50 a share. What is the company's cost of preferred stock, rp? Round your answer to two decimal places.

Percy Motors has a target capital structure of 40% debt and 60% common equity, with no preferred stock. The yield to maturity on the company's outstanding bonds is 9%, and its tax rate is 40%. Percy's CFO estimates that the company's WACC is 13.50%. What is Percy's cost of common equity? Round your answer to two decimal places.
Answer #1

multiply by 1-Tax rate After tax cost of debt=1 7.80% [ 13% (1-40%)]
Answer #2

SOLUTION :


Heuser Company :


After tax cost of debt, rD : 

= Pre tax cost of debt * ( 1 - Tax rate in decimals)

= Yield to maturity in percentage * ( 1 - Tax rate in decimals)

= 13 * ( 1 - 0.40)

= 7.80 % (ANSWER)


Tunny Industries :


Cost of Preferred Stock, rP :

= Dividend / Price of PS

= 6.50 / 71

= 0.0915

= 9.15 % (ANSWER)


Percy Motors :


WACC = wS * rS + wP * rP + wD * rD(1 - T)

=> 13.50 = 0.60 * rS + 0 * rP + 0.40 * 9(1 - 0.40).  (since rD = YTM)

=> 13.50 = 0.60 * rS + 0 + 2.16

=> rS = cost of common stock = (13.50 - 2.16) / 0.60 = 18.90 % (ANSWER).

answered by: Tulsiram Garg